Angle:

I actually need help Jake takes 6 hours to complete a particular job. Suppose he and John begin working on the job together, after 2 hours Jake leaves. If John completes the job himself an hour later, how long would it have taken John to complete the same job if he worked alone?

Angle:

so what I have so far: Jake by himself = 6 hours Some Combination = (2 hours of J+J) + (1 hour of John by himself) want to find John by himself = ? hours

Bob:

the johns and jakes make it confusing, why couldn't it be 2 more different names lol

Angle:

I know right haha how about this A = Jake B = John

Bob:

D = Rate * time Wouldn't we use this formula

Bob:

where d represents number of houses and not distance

Angle:

yeah but I guess it would be a sum like (combined rate)*(time of combined) + (john rate)*(time john took) = 1 whole job

Bob:

im trying to bring nightowl into this lmaooo

Bob:

x = rate of jake y = rate of john 1 = x * 6 would be jake 1 = (x + y) * 2 + y * 1 expanded 1= 6x 1 = 2x + 3y solve for x x = 1/6 substitute into the 2nd equation 1 = 2 (1/6) + 3y 1 = 2/6 + 3y 1 - 2/6 = 3y 1 - 1/3 = 3y 2/3 = 3y y = 2/3 * 1/3 y = 2/9 hours

Bob:

@BlankSpace nightowl's a genius fam lol

Angle:

I just got to that answer myself.... but that means that John is a freaking demon doing a job that takes Jake 6 hours, he does it in less than 14 minutes

Angle:

I was doubting myself for a good few minutes but thank nightowl for me

Bob:

he says he only has an undergraduate math degree lol

Shadow:

John is Barry Allen

Angle:

this is... from my high school math course i.e. high school math content for future teachers

Bob:

lol shadow

Bob:

wait nightowl is telling me otherwise

Bob:

says 2/9 is the rate at which John completes a single house but to find the time... 1 = 2/9 * t 9/2 hours = t

Bob:

4.5 hours

Angle:

that makes more sense x'D

Ultrilliam:

One of these days we will be able to get nightowl aboard the QC ship... x'D

dude:

He exists.